OpenStudy (anonymous):

What is functional equation of \[exp(x+y)\] and \[x,y \in \mathbb{R}\] and conclude from solution \[\sqrt{exp(x)} = exp(\frac{x}{2})\]

OpenStudy (anonymous):

i think functional equation from \[exp(x+y)\] is \[exp(x+y)=exp(x).exp(y)\] but i dont know how to conclude from that \[\sqrt{exp(x)}=exp(\frac{x}{2})\]

OpenStudy (amistre64):

isnt a funciton equation something like: f(x,y,...,k) = 0

OpenStudy (amistre64):

we know that:\[\sqrt{e^x}=(e^x)^{1/2}=e^{x/2}\]

OpenStudy (anonymous):

Nice explaining :)

OpenStudy (anonymous):

hmm maybe, i am not sure but i can write what in my german book is as definition for functional equation of exponential function is like that \[e^{x+y}={e^{x}}.{e^{y}} \forall x, y \in \mathbb{R} \]

OpenStudy (amistre64):

wikipedia is saying your f(x+y) = f(x)f(y) definition is fine as well

OpenStudy (anonymous):

ooh ok, also what would be the functional equation of \[ \exp(x+y)\] ?

OpenStudy (amistre64):

hmm, i dont think i have a good idea on how to process this just yet. exp(x+y) = exp(x) exp(y) if we let x=y then exp(x+x) = exp(x) exp(x) exp(2x) = exp^2(x) is what im thinking of

OpenStudy (anonymous):

ok amistre , thank you very much for nice explanation

OpenStudy (amistre64):

good luck :)